OpenStudy (anonymous):

Context; My son and I have started a business manufacturing and selling his invention. It is a backboard for rescuing people but it has the features of a dolly built in. (www.waukboard.com) I want to show the advantage, of rolling the patient instead of carrying the patient - the Problem - if the patient and the WauK(TM) board weigh 300lbs, and 2 firemen rolling the patient lift the head end to shoulder height and then begin rolling the patient 1 - is the initial dead-lift 75lbs per firemen to get the patient to shoulder height? - 2 - at the angles of 75 deg. and 55 deg how much force is required for each of the fireman as they roll the 300lb patient forward?

OpenStudy (anonymous):

The patient exert 300lbs downward which is at their centre of gravity(at the mid-point when the total lenght is checked)so th force on the firemen shoulder would be less than 75lbs remember that force decreaes as it goes away from the centre of gravity. The second question now, i came up with a formular which is Fsin75=300lbs, if you make F the subject we have F=300/sin75 that goes to angle 55 also. I did this because when you draw the moment diagram and obey one of the rules which says sum of upward forces=sum of download forces..the 75 an 55 degree are both inclined to the horizontal,so when it is being resolved vertically, it's becomes Fsin theta

OpenStudy (anonymous):

Bolatito sin of 75 deg = .9659 therefore the answer for the force required to hold the 300lbs at the head end of the board for the 2 fireman is 300/.9659 = 310.6 lbs this doesn't make any sense garry

OpenStudy (mos1635):

 W=F+N assume normal destribution of weight W*cosθ*L/2=F*cosθ*L F=W/2